The original fault with this set was loss of raster, with the
power supply module going into its “purring” mode. The
cause was found to be a defective horizontal shift choke
(L752), and replacing this restored the sound and picture.
There are still a couple of faults however. First the picture is
of reduced width, the width control 8860 (in the EW
modulator circuit) having no effect at all. The voltage at the
supply end reads 30V instead of 25V. Secondly, although
the 6.8V reference zener diode D975 in the switch -mode
power supply control circuit has been changed (to cure picture size variations) the h.t. supply cannot be set any higher
than 150V (should be 160V). The set h.t. control R976 can
be used to reduce the h.t. below 150V however. Whilst
checking around the circuit I noticed that the EW
modulator diodes D750 and D751 seem to have been running hot, as the board is slightly scorched in the vicinity.
All your problems are likely to be resolved by replacing
the EW modulator diodes, plus the EW modulator output
transistors Q853/852 if necessary. Also check whether the
horizontal shift control R772 is burnt.